OpenStudy (anonymous):

Can someone show me the steps to this differentiation? \[\frac { d }{ dx } { 2 }^{ x }\] How should I start with differentiating this?

sam (.sam.):

2^x

OpenStudy (anonymous):

Take log on both sides, y=2^x

OpenStudy (anonymous):

\[y=2^{x}\] \[lny=xln2\] \[[lny]\prime=\ln2\] \[y'/y=\ln2\] \[y'=\ln2*y'\] \[y'=2^{x}*\ln2\]

sam (.sam.):

d/dx e^x : by d/dx ln(x) Given : d/dx ln(x) = 1/x; Chain Rule; d/dx x = 1. (1) d/dx ln(e^x) = d/dx x = 1 d/dx ln(e^x) = d/du ln(u) d/dx e^x (Set u=e^x) = 1/u d/dx e^x = 1/e^x d/dx e^x = 1 (equation 1) d/dx e^x = e^x

OpenStudy (anonymous):

ohh...Thank you!!!

OpenStudy (anonymous):

sorry, second to the last line should read y' = ln2 * y